SX AO Calibration

New Pupil motors position
BLUE 330 420
YELLOW 371 443
RED 495 402

New calibration 150 modes in OL
MH LGSW IM LGSW REC COMBO LGSW REC COMBO LGSW+BIN4 REC
20170302_183150 20170303_175221 20170303_175400 20170303_175500 20170303_175600

Test new recontrstuctor with 1.0" disturbance. Adsec star to Skip Frame due to big forces.
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Plot
20170303_180505 0.55,0.5,0.5 F 67.4312 85.177 20170303_175600 -1 T F 20170303_180505-20170303_180529_modalplot.png
20170303_180516 0.55,0.5,0.5 T 80.6067 97.0494 20170303_175600 -1 T F
20170303_180529 0.55,0.5,0.5 T 84.3393 133.025 20170303_175600 -1 T F

Old Rec:
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Plot
20170303_185909 0.5,0.45,0.4 T 152.119 79.6799 20161013_153800 -1 F F 20170303_185909-20170303_185956_modalplot.png
20170303_185937 0.5,0.5,0.5 T 217.366 112.713 20161013_153800 -1 F F
20170303_185956 0.5,0.5,0.4 T 147.447 83.7595 20161013_153800 -1 F F
New IM acquisition in OL and CL with different ModalHistory: no good.

Start acquisition of three independent IM. To be finished.

New flat on SX adsec with 400 modes on FLAO: 20170303_flao400modes.sav

150 modes in OL, jitter gain 0.4 (swing arm was oscillating).
MH LGSW IM LGSW REC COMBO LGSW REC
  20170303_215007 20170303_215000 20170303_21550

150 modes in CL 20170303_221100, jitter gain 0.4 (swing arm was oscillating), gain TT=0.0.
MH LGSW IM LGSW REC COMBO LGSW REC COMBO LGSW+BIN4 REC COMBO LGSW+BIN2 REC COMBO LGSW+QC REC
  20170303_220520 20170303_21100 20170303_221200 20170303_221300 20170306_202900 20170306_203000

New Pupil motors position
BLUE 325 425
YELLOW 371 443
RED 500 412

Test new recontrstuctor with 1.0" disturbance.
TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Plot
20170303_225456 0.5,0.4,0.4 T 126.88 105.88 20170303_221300 -1 T F 20170303_225456-20170303_225549_modalplot.png
20170303_225525 0.5,0.4,0.4 T 141.062 102.628 20170303_221300 -1 T F
20170303_225549 0.5,0.4,0.4 F 140.306 107.265 20170303_221300 -1 T F

We try to remove the come aberration with the HowToRemoveComa procedure. New hexapod position:
SX HEXA 5.87 2.18 -0.12 60 193
-- TommasoMazzoni - 14 Mar 2017
Topic revision: r4 - 28 Apr 2017, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback